OpenStudy (anonymous):

Suppose you take out a 1000$ loan for tuition at UMASS on sep. 1st and the loan charges 9% interests. IN december, you win a scholarship with a 2000$ stipend and decide to pay off the loan on January 1st. How much do you need to pay back? HELP PLEASE.

OpenStudy (anonymous):

Is the 9% interest monthly? annually?

OpenStudy (anonymous):

it doesn't say, which is why i'm bit confused on how to start this. :(

OpenStudy (anonymous):

Well, you can probably assume that the interest rate is annually. But then again it depends on if they consider a part of a year to be worth a year or if they do increments... I would say that you would have to pay 1090 back (1000 + (1000*0.09))

OpenStudy (anonymous):

hm. okay thanks though! :)

OpenStudy (anonymous):

sure thing :)

OpenStudy (anonymous):

It seems like 9% is monthly interest, since it mentions the time from 1st sept. to 1st Jan!

OpenStudy (anonymous):

really? So you would use the equation, A=p(1+r/k)^kt?

OpenStudy (anonymous):

Yes, that's how I observe :)

OpenStudy (anonymous):

Well if it's monthly interest, my above answer is wrong D:

OpenStudy (anonymous):

@wombat we just trying to analyze the post! No guarantee right or wrong :P

OpenStudy (anonymous):

Haha im playing pity party here :P

OpenStudy (anonymous):

A = 1,000 ( 1 + .09/12 ) ^ (4/12) = $ 1,002.5

OpenStudy (anonymous):

Can you continue?

OpenStudy (anonymous):

so 1,002.5 is how much they owe?

OpenStudy (anonymous):

The loan amount she owns!

OpenStudy (anonymous):

okay. i see. thanks. one more quick question. how do you find out it if is annually, monthly, or continuosly by looking at the word problem?

OpenStudy (anonymous):

If the question doesn't mention in detail about the time, then it's annually! If it's continously, the question will name it out most of the time!

OpenStudy (anonymous):

Almost all of loan type compounded monthly!

OpenStudy (anonymous):

okay thanks! :D